Status Report

NOAA GOES SATOPS Morning Report 21 Dec 2000

By SpaceRef Editor
December 21, 2000
Filed under ,

SATOPS MORNING REPORT – December 21, 2000

POES

No change in the status of the POES spacecraft.

NOAA-15 rev 13538 / W at 1713Z on December 20:  Commanded spacecraft
to place MIRP in AVHRR synch.  Returned MIRP to internal synch on
rev 13540.

POES operations were nominal over the past 24 hours.

DMSP

No change in the status of the DMSP spacecraft.

At 1610Z and 1626Z on December 20:  Nominal pre-launch 10K loop-back
testing through Thule Tracking Station, Greenland.

DMSP operations were nominal over the past 24 hours.

GOES

No change in the status of the GOES spacecraft.

At 1758Z on December 20:  GOES-10 image break due to loss of signal
synch at CDA.

GOES operations were nominal over the past 24 hours.

 

Questions or comments to [email protected]

SpaceRef staff editor.